OpenStudy (anonymous):

Jakob is purchasing a condominium for $347,500 and has financed 90% of it. He has purchased 2 discount points, has to pay a 1.75% origination fee and a brokerage fee of $100 plus 0.25% of the loan. What are his total fees that he must pay?

jimthompson5910 (jim_thompson5910):

what were your answer choices again?

OpenStudy (anonymous):

@ jim_thompson5910 $12,610.01 $13,031.25 $13,218.13 $13,913.13

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

if he bought it for "$347,500 and has financed 90% of it." then how much is being loaned to him?

OpenStudy (anonymous):

it is not mentioned :/

jimthompson5910 (jim_thompson5910):

well what's 90% of the price of the home

OpenStudy (anonymous):

312,750

jimthompson5910 (jim_thompson5910):

so that's the amount being financed (ie loaned out)

jimthompson5910 (jim_thompson5910):

He has purchased 2 discount points what does that mean?

jimthompson5910 (jim_thompson5910):

this is a very important home buying term/concept

OpenStudy (anonymous):

2 discount points is basically like 2% of your loan amount.

jimthompson5910 (jim_thompson5910):

what about 2% of your loan

OpenStudy (anonymous):

then that is 6,255

jimthompson5910 (jim_thompson5910):

so that's how much it costs to purchase 2 discount points

jimthompson5910 (jim_thompson5910):

then he " has to pay a 1.75% origination fee " so 1.75% of price of home = ??

OpenStudy (anonymous):

5,254.20?

jimthompson5910 (jim_thompson5910):

312750*0.0175 = 5,473.125

jimthompson5910 (jim_thompson5910):

so that rounds to 5473.13

jimthompson5910 (jim_thompson5910):

that's the origination fee

jimthompson5910 (jim_thompson5910):

then you have to add on " a brokerage fee of $100 plus 0.25% of the loan"

jimthompson5910 (jim_thompson5910):

what do you get for that part

OpenStudy (anonymous):

5,573.13

OpenStudy (anonymous):

hold on I did something wrong

jimthompson5910 (jim_thompson5910):

you might have done 25% it's 0.25%

OpenStudy (anonymous):

100 plus 5,473.13=5,573.13+0.25=5,587.06

jimthompson5910 (jim_thompson5910):

that's too big, the brokerage fee is much smaller

jimthompson5910 (jim_thompson5910):

again, it's NOT 25% it's 0.25%

jimthompson5910 (jim_thompson5910):

it should be this 100+0.0025*312750 = 881.875 so the brokerage fee is $881.88

jimthompson5910 (jim_thompson5910):

now add up the discount points, the origination fee, and the brokerage fee to get 6255+5473.13+881.88 = 12,610.01

jimthompson5910 (jim_thompson5910):

keep in mind that this does NOT include the 10% down payment

OpenStudy (anonymous):

Awesome! Thank you. May you help me with one more please?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Shayla is financing a house for $126,750. She has to pay $450 plus 1.15% for a brokerage fee. How much are the mortgage brokerage fees? $2,351.25 $1,907.63 $1,457.63 $595.76 for this one I put $1,907.63***

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

nice work

OpenStudy (anonymous):

Thank you so much for your help:)!

jimthompson5910 (jim_thompson5910):

you're welcome